# Existence and Uniqueness of Solutions for Steady Triple-Deck Equations

### v1

@@ 2. Preliminaries
Lemma 3.1.
Fix $\alpha>0$. We assume that $h_{e}\in H_{x}^{\frac{\alpha}{2}}L_{y}^{2}$. Then, there exists a unique solution $f_{e}\in X_{0}\cap X_{\alpha}$ to the problem (3.1). Moreover, there holds
@@
where $C$ is a constant independent of $\alpha,M$, and $h_{e}$.
@@
Step 1. We give the explicit form of the solution. Taking the Fourier transform in variable $x$ into (3.1), we obtain
@@
Using the Green-function method, we find the unique solution $\hat{f}_{e}$ of (3.2), which has an explicit formula as following:
@@
with
(3.4) $G(y,z;\xi)=\pi s_{M}\left\{\begin{aligned} \mu_{M}(\xi,y)\mu_{0}(\xi,z),&\quad 0\leq z<y,\\ \mu_{M}(\xi,z)\mu_{0}(\xi,y),&\quad y<z\leq M.\end{aligned}\right$.
@@
Step 2. We derive the low-frequency estimates of $f_{e}$. Since $\frac{Ai^{\prime}(0)}{Bi^{\prime}(0)}=-\frac{1}{\sqrt{3}}$, we have
(3.5) $t(z):=\frac{Ai(z)}{Bi(z)}-\frac{Ai^{\prime}(0)}{Bi^{\prime}(0)}=\frac{Ai(z)}{Bi(z)}+\frac{1}{\sqrt{3}}=\frac{\sqrt{3}Ai(z)+Bi(z)}{\sqrt{3}Bi(z)}.$
@@
Then, $t(z)$ is a holomorphic and non-vanishing function on $C_{\frac{\pi}{3}}$. Moreover, for any $\delta\in(0,\frac{\pi}{3})$, there holds
@@
Since $s_{M}=\frac{1}{t((i\xi)^{\frac{1}{3}}M)}$, we have
@@
Then for $|\xi|\leq M^{-3}$, we obtain
(3.9) $|\mu_{M}(\xi,y)|=\bigg{|}Ai((i\xi)^{\frac{1}{3}}y)-\frac{Ai((i\xi)^{\frac{1}{3}}M)}{Bi((i\xi)^{\frac{1}{3}}M)}Bi((i\xi)^{\frac{1}{3}}y)\bigg{|}$
$\qquad\qquad\ \leq C|\xi|^{\frac{1}{3}}(M-y),\quad\forall\ y\in[0,M],$
$|\mu_{0}(\xi,y)|=\bigg{|}Ai((i\xi)^{\frac{1}{3}}y)+\frac{1}{\sqrt{3}}Bi((i\xi)^{\frac{1}{3}}y)\bigg{|}\leq C,\quad\forall\ y\in[0,M].$
@@
(3.10) $|\hat{f}_{e}(\xi,y)|\leq$ $\ C|s_{M}|\cdot|\xi|^{-\frac{1}{3}}\bigg{\{}\big{|}\mu_{M}(\xi,y)\big{|}\cdot\int_{0}^{y}\big{|}\mu_{0}(\xi,z)\hat{h}_{e}(\xi,z)\big{|}dz$
@@
Hence, we have
@@
And for $|\xi|\geq M^{-3}$, we have
@@
(3.14) $|\hat{f}_{e}(\xi,y)|\leq C|s_{M}|\cdot|\xi|^{-\frac{1}{3}}\bigg{\{}\big{|}\mu_{M}(\xi,y)\big{|}\cdot\int_{0}^{y}\big{|}\mu_{0}(\xi,z)\hat{h}_{e}(\xi,z)\big{|}dz$
@@
Hence, we have
@@
Step 3. We derive the high-frequency estimates of $f_{e}$. For any $\bar{\alpha}\geq 0$, we set
@@
Multiplying the equation (3.2) by $\xi^{1+\bar{\alpha}}y\hat{f}_{e}$, we have
(3.18) $i|\xi|^{2+\bar{\alpha}}{\big{\|}y\hat{f}_{e}\big{\|}}_{L_{y}^{2}}^{2}+\xi^{1+\bar{\alpha}}{\big{\|}y^{\frac{1}{2}}\partial_{y}\hat{f}_{e}\big{\|}}_{L_{y}^{2}}^{2}+\xi^{1+\bar{\alpha}}\langle\partial_{y}\hat{f}_{e},\hat{f}_{e}\rangle_{L_{y}^{2}}=\langle\hat{h}_{e},y\xi^{1+\bar{\alpha}}\hat{f}_{e}\rangle_{L_{y}^{2}}.$
Here we use a fact that $\hat{f}_{e}\big{|}_{y=M}=0$. Set
@@
$\qquad\leq\big{|}-\langle\xi^{\frac{1}{3}+\frac{\bar{\alpha}}{2}}\partial_{y}\hat{f}_{e},\xi^{\frac{2}{3}+\frac{\bar{\alpha}}{2}}\hat{f}_{e}\rangle_{L_{y}^{2}}+\langle\xi^{\frac{\bar{\alpha}}{2}}\hat{h}_{e},y\xi^{1+\frac{\bar{\alpha}}{2}}\hat{f}_{e}\rangle_{L_{y}^{2}}\big{|}$
$\qquad\leq C\bigg{(}|\xi|^{\frac{1}{3}+\frac{\bar{\alpha}}{2}}{\big{\|}\partial_{y}\hat{f}_{e}\big{\|}}_{L_{y}^{2}}+|\xi|^{\frac{\bar{\alpha}}{2}}{\big{\|}\hat{h}_{e}\big{\|}}_{L_{y}^{2}}\bigg{)}B_{0}$
$\qquad\leq CB_{0}(B_{2}+|\xi|^{\frac{\bar{\alpha}}{2}}{\big{\|}\hat{h}_{e}\big{\|}}_{L_{y}^{2}}).$
@@
Then we get
(3.22) $|\xi|^{\frac{5}{3}+\bar{\alpha}}{\big{\|}y^{\frac{1}{2}}\hat{f}_{e}\big{\|}}_{L_{y}^{2}}^{2}+|\xi|^{\frac{2}{3}+\bar{\alpha}}{\big{\|}\partial_{y}\hat{f}_{e}\big{\|}}_{L_{y}^{2}}^{2}\leq{\big{\|}|\xi|^{\frac{2}{3}+\frac{\bar{\alpha}}{2}}\hat{f}_{e}\big{\|}}_{L_{y}^{2}}{\big{\|}|\xi|^{\frac{\bar{\alpha}}{2}}\hat{h}_{e}\big{\|}}_{L_{y}^{2}}$
$\qquad\leq CB_{0}{\big{\|}|\xi|^{\frac{\bar{\alpha}}{2}}\hat{h}_{e}\big{\|}}_{L_{y}^{2}}.$
Thus, by (3.19), we obtain
@@
On the other hand, it is straight to obtain
@@
$\qquad=|\xi|^{\frac{4}{3}+\bar{\alpha}}M\hat{f}_{e}^{2}\big{|}_{y=M}-|\xi|^{\frac{4}{3}+\bar{\alpha}}\int_{0}^{M}y(\bar{\hat{f}}_{e}\partial_{y}\hat{f}_{e}+\hat{f}_{e}\partial_{y}\bar{\hat{f}}_{e})dy$
$\qquad\leq 2{\big{\|}|\xi|^{1+\frac{\bar{\alpha}}{2}}y\hat{f}_{e}\big{\|}}_{L_{y}^{2}}{\big{\|}|\xi|^{\frac{1}{3}+\frac{\bar{\alpha}}{2}}\partial_{y}\hat{f}_{e}\big{\|}}_{L_{y}^{2}}$
$\qquad\leq CB_{0}B_{2}.$
By (3.17), (3.19), (3.20), (3.23), and (3.24), we have
@@
Applying Plancherel Theorem, we obtain
@@
Here $C$ is independent of $\alpha,M$, and $f_{e}$.
@@
The Airy functions $Ai(z),Bi(z)$ admit decompositions [20, Chapter 9] as follows.
(3.29) $Ai(z)=\frac{\sqrt{z}}{3}\big{(}I_{-\frac{1}{3}}(\frac{2}{3}z^{\frac{3}{2}})-I_{\frac{1}{3}}(\frac{2}{3}z^{\frac{3}{2}})\big{)},\quad z\in{\mathbb{C}},$
@@
For $\nu>-1$, the Bessel function $J_{\nu}(z)$ has only real zeros in $\mathbb{C}$} [31, Page 482]. Then by (3.30), the zeros of ${\mathcal{C}}(z)$ in the sector $\big{\{}z\in{\mathbb{C}}|\ -\frac{2\pi}{3}\leq\text{Arg}(z)\leq\frac{\pi}{3}\big{\}}$ lie only on rays $\big{\{}z\in{\mathbb{C}}|\ \text{Arg}(z)=-\frac{\pi}{3}\ \text{or}\ \frac{\pi}{3}\big{\}}$. Note that ${\mathcal{C}}(z)$ is entire and ${\mathcal{C}}(z)={\mathcal{C}}(ze^{\frac{2\pi i}{3}})$ for any $z\in{\mathbb{C}}$, then we have that all zeros of ${\mathcal{C}}(z)$ in $\mathbb{C}$} belong to $\big{\{}z\in{\mathbb{C}}|\ \text{Arg}(z)=\frac{\pi}{3}+\frac{2k\pi}{3},\ k\in{\mathbb{Z}}\big{\}}$.
By homogenization of (1.15), it is straight to obtain the linearized problem:
(3.31) $\left\{\begin{aligned} &y\partial_{x}f-\partial_{y}^{2}f=0,\quad(x,y)\in\Omega_{M},\\ &\partial_{y}f\big{|}_{y=0}=\partial_{x}|\partial_{x}|A,\quad I_{M}[f]=A+F,\quad x\in{\mathbb{R}}.\end{aligned}\right$.
In order to solve the problem (3.31), we first omit the integral condition.
Lemma 3.2.
Let $\bar{\alpha}\geq 0$ and $g=\partial_{x}|\partial_{x}|A\in H^{\frac{1}{6}+\frac{\bar{\alpha}}{2}}\cap\dot{H}^{-1}$. Then, there exists a solution $\omega_{b}\in X_{\bar{\alpha}}$ of the boundary-value problem:
(3.32) $\left\{\begin{aligned} &y\partial_{x}\omega_{b}-\partial_{y}^{2}\omega_{b}=0,\quad(x,y)\in\Omega_{M},\\ &\partial_{y}\omega_{b}\big{|}_{y=0}=g,\quad x\in{\mathbb{R}},\end{aligned}\right$.
with
@@
To solve the boundary-value problem (3.32), we first take the Fourier transform in variable $x$. Then we obtain
@@
The general solution to this BVP (3.34) is
@@
Taking $C(\xi)=0$, we obtain a solution $\omega_{b}$ with the explicit formula as following:
@@
$\qquad\leq C{\big{\|}{\big{\|}Ai((i\xi)^{\frac{1}{3}}y)(|\xi|^{\frac{1}{3}}y)^{\alpha^{\prime}}\big{\|}}_{L_{y}^{2}}\cdot|\xi|^{\frac{5}{3}}|\hat{A}(\xi)|\big{\|}}_{L_{\xi}^{2}}$
$\qquad\leq C{\big{\|}|\xi|^{\frac{3}{2}}\hat{A}(\xi)\big{\|}}_{L_{\xi}^{2}}=C{\big{\|}|\partial_{x}|^{\frac{3}{2}}A\big{\|}}_{L^{2}({\mathbb{R}})}$
$\qquad\leq C{\big{\|}|\partial_{x}|^{\frac{5}{6}}A\big{\|}}_{H^{\frac{2}{3}}({\mathbb{R}})}.$
Since $g\in H^{\frac{1}{6}+\frac{\alpha}{2}}({\mathbb{R}})$, we obtain
(3.38) ${\big{\|}f\big{\|}}_{Y_{\bar{\alpha}}}=$ $\ {\big{\|}y|\partial_{x}|^{\frac{\bar{\alpha}}{2}}\partial_{x}\omega_{b}\big{\|}}_{L_{x,y}^{2}}+{\big{\|}|\partial_{x}|^{\frac{2}{3}+\frac{\bar{\alpha}}{2}}f\big{\|}}_{L_{x,y}^{2}}+{\big{\|}|\partial_{x}|^{\frac{1+\bar{\alpha}}{2}}y^{\frac{1}{2}}\partial_{y}f\big{\|}}_{L_{x,y}^{2}}$
$\ \ +{\big{\|}|\partial_{x}|^{\frac{1}{3}+\frac{\bar{\alpha}}{2}}\partial_{y}f\big{\|}}_{L_{x,y}^{2}}+{\big{\|}|\partial_{x}|^{\frac{\bar{\alpha}}{2}}\partial_{y}^{2}\omega_{b}\big{\|}}_{L_{x,y}^{2}}$
@@
It follows $\omega_{b}\in X_{\bar{\alpha}}$ from (3.37) and (3.38). In particular, we have
(3.39) ${\big{\|}\omega_{b}\big{\|}}_{X_{\bar{\alpha}}}\leq C{\big{\|}|\partial_{x}|^{\frac{5}{6}+\frac{\bar{\alpha}}{2}}A\big{\|}}_{H^{\frac{4}{3}}({\mathbb{R}})}.$
Now, we are in the position to prove Theorem 1.1. Denote $(\omega_{0},u_{0},v_{0},A_{0})$ be the solution to
@@
We hope that we take (3.36) as a solution to (3.40), that is,
@@
(3.44) $p_{M}(\xi)=$ $\ \int_{0}^{(i\xi)^{\frac{1}{3}}M}Ai(z)dz.$
Indeed, the integral is well-defined since $Ai(z)$ is analytic. We set
@@
To ensure that the formulas (3.46) and (3.47) are well-defined, we need $m(\xi)\not=0$ for any $\xi\in{\mathbb{R}}$. Then, we give an estimate of $m(\xi)$. We define
@@
Note that $0\leq p_{0}<\infty$, then we obtain
@@
(3.48) $m(\xi)\geq\frac{1}{4},\quad\forall\ |\xi|\leq a_{0}.$
@@
(3.49) $\int_{0}^{\infty}Ai(z)dz=\frac{1}{3}.$
As $M\to\infty$, the multiplier $m(\xi)$ converges to $\frac{1}{3}$ uniformly on $a_{0},\infty)$. By calculations, we have
@@
Then there exists a large enough constant $M_{0}>0$, when $M\geq M_{0}$, by (3.48) and (3.50), we also obtain
@@
It concludes that $\hat{A}_{0}$ and $\hat{\omega}_{0}$ are well-defined. Moreover, by Fourier multiplier theorem, there exists a constant $C>0$ such that
(3.52) ${\big{\|}m(D)^{-1}f\big{\|}}_{H^{\frac{4}{3}}({\mathbb{R}})}\leq C{\big{\|}f\big{\|}}_{L^{2}({\mathbb{R}})},\quad\forall\ f\in L^{2}({\mathbb{R}}).$
Using Lemma 3.2 and the estimate (3.52), for any $\bar{\alpha}\geq 0$, we obtain the $X_{\bar{\alpha}}$-estimate of $\omega_{0}$ as following:
(3.53) ${\big{\|}\omega_{0}\big{\|}}_{X_{\bar{\alpha}}}\leq C{\big{\|}|\partial_{x}|^{\frac{5}{6}+\frac{\bar{\alpha}}{2}}A_{0}\big{\|}}_{H^{\frac{4}{3}}({\mathbb{R}})}\leq C{\big{\|}|\partial_{x}|^{\frac{5}{6}+\frac{\bar{\alpha}}{2}}F\big{\|}}_{L^{2}({\mathbb{R}})}\leq C{\big{\|}|\partial_{x}|^{\frac{5}{6}}F\big{\|}}_{\dot{H}^{\frac{\bar{\alpha}}{2}}({\mathbb{R}})}.$
For the bound of displacement $A_{0}$ and the velocity $u_{0}(x)$, we obtain
(3.54) ${\big{\|}u_{0}\big{\|}}_{L^{\infty}(\Omega_{M})}+{\big{\|}A_{0}\big{\|}}_{L^{\infty}({\mathbb{R}})}\leq{\big{\|}\omega_{0}\big{\|}}_{L_{x}^{\infty}L_{y}^{1}}\leq{\big{\|}\omega_{0}\big{\|}}_{L_{y}^{1}L_{x}^{\infty}}\leq{\big{\|}\omega_{0}\big{\|}}_{X_{0}}.$
It completes the proof. ∎

### v2

@@ 3. The solvability of the linear system
Proposition 3.1.
Assume that $h_{e}\in H_{x}^{\frac{\alpha}{2}}L_{y}^{2}$ for some $\alpha>0$. Then there exists a unique solution $f_{e}\in X_{0}\cap X_{\alpha}$ to the problem (3.3). Moreover, there holds that
@@
where $C$ is a constant independent of $\alpha,M$.
@@
Step 1. The explicit formula of the solution.
Taking the Fourier transform in variable $x$ into (3.3), we obtain
@@
Using the Green-function method, we find the unique solution $\hat{f}_{e}$ of (3.4), which has an explicit formula
@@
with the definition of $G$ given by
(3.6) $G(y,z;\xi)=\pi s_{M}\left\{\begin{aligned} \mu_{M}(\xi,y)\mu_{0}(\xi,z),&\quad 0\leq z<y,\\ \mu_{M}(\xi,z)\mu_{0}(\xi,y),&\quad y<z\leq M,\end{aligned}\right$.
@@
Step 2. The low-order estimates of $f_{e}$.
To this end, we first claim that $s_{M}$ is bounded. Since the ratio $\frac{Ai^{\prime}(z)}{Bi^{\prime}(z)}$ of the Airy functions at $z=0$ is $\frac{1}{\sqrt{3}}$, we then define
$t(z):=\frac{Ai(z)}{Bi(z)}-\frac{Ai^{\prime}(0)}{Bi^{\prime}(0)}=\frac{Ai(z)}{Bi(z)}+\frac{1}{\sqrt{3}}=\frac{\sqrt{3}Ai(z)+Bi(z)}{\sqrt{3}Bi(z)}.$
@@
Then, $t(z)$ is a holomorphic and non-vanishing function on $C_{\frac{\pi}{3}}$. Moreover, for any $\delta\in(0,\frac{\pi}{3})$, the Airy functions $Ai,Bi$ satisfy the following asymptotic expansions(see [20, Chapter 9]):
@@
Thanks to $s_{M}=\bigg(\frac{Ai((i\xi)^{\frac{1}{3}}M)}{Bi((i\xi)^{\frac{1}{3}}M)}-\frac{Ai^{\prime}(0)}{Bi^{\prime}(0)}\bigg)^{-1}=\frac{1}{t((i\xi)^{\frac{1}{3}}M)}$, we then obtain
@@
To derive a low-order estimate of $f_{e}$, we decompose the frequency space into two regimes. For the low-frequency regime where $|\xi|\leq M^{-3}$, we observe that $|(i\xi)^{\frac{1}{3}}y|\leq 1$ for all $y\in[0,M]$. By the continuity of $Ai$ and $Bi$ on $\mathbb{C}$}, we have
(3.10) $|\mu_{0}(\xi,y)|=\big|Ai((i\xi)^{\frac{1}{3}}y)+\frac{1}{\sqrt{3}}Bi((i\xi)^{\frac{1}{3}}y)\big|\leq C,\quad\forall\ y\in[0,M].$
Next, we compute
(3.11) $\mu_{M}(\xi,y)=Ai((i\xi)^{\frac{1}{3}}y)-\frac{Ai((i\xi)^{\frac{1}{3}}M)}{Bi((i\xi)^{\frac{1}{3}}M)}Bi((i\xi)^{\frac{1}{3}}y)$
$\qquad\quad\ \ =Ai((i\xi)^{\frac{1}{3}}y)-\frac{Ai\big((i\xi)^{\frac{1}{3}}(M-y)+(i\xi)^{\frac{1}{3}}y\big)}{Bi\big((i\xi)^{\frac{1}{3}}(M-y)+(i\xi)^{\frac{1}{3}}y\big)}Bi((i\xi)^{\frac{1}{3}}y).$
Since $Ai((i\xi)^{\frac{1}{3}}y)-\frac{Ai((i\xi)^{\frac{1}{3}}y)}{Bi((i\xi)^{\frac{1}{3}}y)}Bi((i\xi)^{\frac{1}{3}}y)=0$ and $Bi(z)\not=0$ for ${\rm Arg}(z)\in(-\frac{\pi}{3},\frac{\pi}{3})$, an application of the Mean Value Theorem to (3.11) yields
(3.12) $|\mu_{M}(\xi,y)|\leq C|\xi|^{\frac{1}{3}}(M-y),\quad\forall\ y\in[0,M].$
@@
(3.13) $|\hat{f}_{e}(\xi,y)|=$ $\ \bigg|\pi s_{M}(i\xi)^{-\frac{1}{3}}\bigg\{\mu_{M}(\xi,y)\cdot\int_{0}^{y}\mu_{0}(\xi,z)\hat{h}_{e}(\xi,z)dz$
$\qquad\qquad\qquad\quad+\mu_{0}(\xi,y)\cdot\int_{y}^{M}\mu_{M}(\xi,z)\hat{h}_{e}(\xi,z)dz\bigg\}\bigg|$
$\leq$ $\ C|s_{M}|\cdot|\xi|^{-\frac{1}{3}}\bigg\{\big|\mu_{M}(\xi,y)\big|\cdot\int_{0}^{y}\big|\mu_{0}(\xi,z)\hat{h}_{e}(\xi,z)\big|dz$
@@
Hence, we obtain a point-wise estimate of $\hat{f}_{e}$ for the low frequency as follows
@@
For the high frequency $|\xi|\geq M^{-3}$, by the asymptotic behavior (3.7) of Airy functions, we have
@@
$|\hat{f}_{e}(\xi,y)|\leq C|s_{M}|\cdot|\xi|^{-\frac{1}{3}}\bigg\{\big|\mu_{M}(\xi,y)\big|\cdot\int_{0}^{y}\big|\mu_{0}(\xi,z)\hat{h}_{e}(\xi,z)\big|dz$
@@
Hence, we obtain a pointwise estimate of $\hat{f}_{e}$ for the high frequency as follows
@@
Step 3. The high-order estimates of $f_{e}$.
For $\bar{\alpha}\geq 0$, we set
@@
Multiplying the equation (3.4) by $\xi^{1+\bar{\alpha}}y\hat{f}_{e}$, we obtain
$i|\xi|^{2+\bar{\alpha}}{\big\|y\hat{f}_{e}\big\|}_{L_{y}^{2}}^{2}+\xi^{1+\bar{\alpha}}{\big\|y^{\frac{1}{2}}\partial_{y}\hat{f}_{e}\big\|}_{L_{y}^{2}}^{2}+\xi^{1+\bar{\alpha}}\langle\partial_{y}\hat{f}_{e},\hat{f}_{e}\rangle_{L_{y}^{2}}=\langle\hat{h}_{e},y\xi^{1+\bar{\alpha}}\hat{f}_{e}\rangle_{L_{y}^{2}}.$
Here we used the fact that $\hat{f}_{e}\big|_{y=M}=0$. Set
@@
$\leq\big|-\langle\xi^{\frac{1}{3}+\frac{\bar{\alpha}}{2}}\partial_{y}\hat{f}_{e},\xi^{\frac{2}{3}+\frac{\bar{\alpha}}{2}}\hat{f}_{e}\rangle_{L_{y}^{2}}+\langle\xi^{\frac{\bar{\alpha}}{2}}\hat{h}_{e},y\xi^{1+\frac{\bar{\alpha}}{2}}\hat{f}_{e}\rangle_{L_{y}^{2}}\big|$
$\leq C\bigg(|\xi|^{\frac{1}{3}+\frac{\bar{\alpha}}{2}}{\big\|\partial_{y}\hat{f}_{e}\big\|}_{L_{y}^{2}}+|\xi|^{\frac{\bar{\alpha}}{2}}{\big\|\hat{h}_{e}\big\|}_{L_{y}^{2}}\bigg)B_{0}$
$\leq CB_{0}(B_{2}+|\xi|^{\frac{\bar{\alpha}}{2}}{\big\|\hat{h}_{e}\big\|}_{L_{y}^{2}}).$
@@
Then we have
$|\xi|^{\frac{5}{3}+\bar{\alpha}}{\big\|y^{\frac{1}{2}}\hat{f}_{e}\big\|}_{L_{y}^{2}}^{2}+|\xi|^{\frac{2}{3}+\bar{\alpha}}{\big\|\partial_{y}\hat{f}_{e}\big\|}_{L_{y}^{2}}^{2}$ $\leq{\big\||\xi|^{\frac{2}{3}+\frac{\bar{\alpha}}{2}}\hat{f}_{e}\big\|}_{L_{y}^{2}}{\big\||\xi|^{\frac{\bar{\alpha}}{2}}\hat{h}_{e}\big\|}_{L_{y}^{2}}$
$\leq CB_{0}{\big\||\xi|^{\frac{\bar{\alpha}}{2}}\hat{h}_{e}\big\|}_{L_{y}^{2}}.$
Thus, we get by (3.20) that
@@
On the other hand, we have
@@
$=|\xi|^{\frac{4}{3}+\bar{\alpha}}M\hat{f}_{e}^{2}\big|_{y=M}-|\xi|^{\frac{4}{3}+\bar{\alpha}}\int_{0}^{M}y(\bar{\hat{f}}_{e}\partial_{y}\hat{f}_{e}+\hat{f}_{e}\partial_{y}\bar{\hat{f}}_{e})dy$
$\leq 2{\big\||\xi|^{1+\frac{\bar{\alpha}}{2}}y\hat{f}_{e}\big\|}_{L_{y}^{2}}{\big\||\xi|^{\frac{1}{3}+\frac{\bar{\alpha}}{2}}\partial_{y}\hat{f}_{e}\big\|}_{L_{y}^{2}}$
$\leq CB_{0}B_{2}.$
By (3.19), (3.20), (3.21), (3.23) and (3.24), we have
@@
Applying Plancherel’s theorem, we obtain
@@
Here $C$ is independent of $\alpha,M$, and $f_{e}$. ∎
@@
The Airy functions $Ai(z),Bi(z)$ admit decompositions [20, Chapter 9] as follows
$Ai(z)=\frac{\sqrt{z}}{3}\big(I_{-\frac{1}{3}}(\frac{2}{3}z^{\frac{3}{2}})-I_{\frac{1}{3}}(\frac{2}{3}z^{\frac{3}{2}})\big),\quad z\in{\mathbb{C}},$
@@
For $\nu>-1$, the Bessel function $J_{\nu}(z)$ has only real zeros in $\mathbb{C}$} [31, Page 482]. Then by (3.29), the zeros of ${\mathcal{C}}(z)$ in the sector $\big\{z\in{\mathbb{C}}|\ -\frac{2\pi}{3}\leq\text{Arg}(z)\leq\frac{\pi}{3}\big\}$ lie only on rays $\big\{z\in{\mathbb{C}}|\ \text{Arg}(z)=-\frac{\pi}{3}\ \text{or}\ \frac{\pi}{3}\big\}$. Notice that ${\mathcal{C}}(z)$ is entire and ${\mathcal{C}}(z)={\mathcal{C}}(ze^{\frac{2\pi i}{3}})$ for any $z\in{\mathbb{C}}$, then we have that all zeros of ${\mathcal{C}}(z)$ in $\mathbb{C}$} belong to $\big\{z\in{\mathbb{C}}|\ \text{Arg}(z)=\frac{\pi}{3}+\frac{2k\pi}{3},\ k\in{\mathbb{Z}}\big\}$.
We next solve the following homogeneous problem
(3.30) $\left\{\begin{aligned} &y\partial_{x}\omega_{b}-\partial_{y}^{2}\omega_{b}=0,\quad(x,y)\in\Omega_{M},\\ &\partial_{y}\omega_{b}\big|_{y=0}=\partial_{x}|\partial_{x}|A,\quad x\in{\mathbb{R}}.\end{aligned}\right$.
Proposition 3.2.
Let $\bar{\alpha}\geq 0$ and $g=\partial_{x}|\partial_{x}|A\in H^{\frac{1}{6}+\frac{\bar{\alpha}}{2}}\cap\dot{H}^{-1}$. Then there exists a solution $\omega_{b}\in X_{\bar{\alpha}}$ to the boundary-value problem (3.30) with the estimate
@@
To solve the boundary-value problem (3.30), we take the Fourier transform in variable $x$. Then we have
@@
The general solution to the BVP (3.32) is as follows
@@
Taking $C(\xi)=0$, we obtain a solution $\omega_{b}$ with the explicit formula
@@
$\leq C{\big\|{\big\|Ai((i\xi)^{\frac{1}{3}}y)(|\xi|^{\frac{1}{3}}y)^{\alpha^{\prime}}\big\|}_{L_{y}^{2}}\cdot|\xi|^{\frac{5}{3}}|\hat{A}(\xi)|\big\|}_{L_{\xi}^{2}}$
$\leq C{\big\||\xi|^{\frac{3}{2}}\hat{A}(\xi)\big\|}_{L_{\xi}^{2}}=C{\big\||\partial_{x}|^{\frac{3}{2}}A\big\|}_{L^{2}({\mathbb{R}})}$
$\leq C{\big\||\partial_{x}|^{\frac{5}{6}}A\big\|}_{H^{\frac{2}{3}}({\mathbb{R}})}.$
As $g\in H^{\frac{1}{6}+\frac{\alpha}{2}}({\mathbb{R}})$, we obtain
(3.36) ${\big\|\omega_{b}\big\|}_{Y_{\bar{\alpha}}}=$ $\ {\big\|y|\partial_{x}|^{\frac{\bar{\alpha}}{2}}\partial_{x}\omega_{b}\big\|}_{L_{x,y}^{2}}+{\big\||\partial_{x}|^{\frac{2}{3}+\frac{\bar{\alpha}}{2}}\omega_{b}\big\|}_{L_{x,y}^{2}}+{\big\||\partial_{x}|^{\frac{1+\bar{\alpha}}{2}}y^{\frac{1}{2}}\partial_{y}\omega_{b}\big\|}_{L_{x,y}^{2}}$
$\ \ +{\big\||\partial_{x}|^{\frac{1}{3}+\frac{\bar{\alpha}}{2}}\partial_{y}\omega_{b}\big\|}_{L_{x,y}^{2}}+{\big\||\partial_{x}|^{\frac{\bar{\alpha}}{2}}\partial_{y}^{2}\omega_{b}\big\|}_{L_{x,y}^{2}}$
@@
It follows from (3.35) and (3.36) that $\omega_{b}\in X_{\bar{\alpha}}$ with
${\big\|\omega_{b}\big\|}_{X_{\bar{\alpha}}}\leq C{\big\||\partial_{x}|^{\frac{5}{6}+\frac{\bar{\alpha}}{2}}A\big\|}_{H^{\frac{4}{3}}({\mathbb{R}})}.$
Now we are in a position to address the system (3.1), i.e., complete the proof of Theorem 1.1. Let $(\omega_{0},u_{0},v_{0},A_{0})$ be the solution to the linear system
@@
We take (3.34) as a solution to (3.37), that is,
@@
$p_{M}(\xi)=$ $\ \int_{0}^{(i\xi)^{\frac{1}{3}}M}Ai(z)dz.$
Indeed, the integral is well-defined, since $Ai(z)$ is analytic. Let
@@
To ensure that the formulas (3.42) and (3.43) are well-defined, we need to prove $m(\xi)\not=0$ for any $\xi\in{\mathbb{R}}$. We define
@@
Noting that $0\leq p_{0}<\infty$, we obtain
@@
(3.44) $m(\xi)\geq\frac{1}{4}\quad\forall\ |\xi|\leq a_{0}.$
@@
(3.45) $\int_{0}^{\infty}Ai(z)dz=\frac{1}{3},$
the multiplier $m(\xi)$ converges to $\frac{1}{3}$ uniformly on $a_{0},\infty)$ as $M\to\infty$. A direct calculation yields
@@
Then there exists a large enough constant $M_{0}>0$ such that for $M\geq M_{0}$, we get by (3.44) and (3.46) that
@@
Thus, $\hat{A}_{0}$ and $\hat{\omega}_{0}$ are well-defined. Moreover, there exists a constant $C>0$ such that
(3.48) ${\big\|m(D)^{-1}f\big\|}_{H^{\frac{4}{3}}({\mathbb{R}})}\leq C{\big\|f\big\|}_{L^{2}({\mathbb{R}})}.$
Using Proposition 3.2 and (3.48), for any $\bar{\alpha}\geq 0$, we obtain the $X_{\bar{\alpha}}$-estimate of $\omega_{0}$ as follows
${\big\|\omega_{0}\big\|}_{X_{\bar{\alpha}}}\leq C{\big\||\partial_{x}|^{\frac{5}{6}+\frac{\bar{\alpha}}{2}}A_{0}\big\|}_{H^{\frac{4}{3}}({\mathbb{R}})}\leq C{\big\||\partial_{x}|^{\frac{5}{6}+\frac{\bar{\alpha}}{2}}F\big\|}_{L^{2}({\mathbb{R}})}\leq C{\big\||\partial_{x}|^{\frac{5}{6}}F\big\|}_{\dot{H}^{\frac{\bar{\alpha}}{2}}({\mathbb{R}})}.$
For the bound of displacement $A_{0}$ and the velocity $u_{0}(x)$, we have
${\big\|u_{0}\big\|}_{L^{\infty}(\Omega_{M})}+{\big\|A_{0}\big\|}_{L^{\infty}({\mathbb{R}})}\leq{\big\|\omega_{0}\big\|}_{L_{x}^{\infty}L_{y}^{1}}\leq{\big\|\omega_{0}\big\|}_{L_{y}^{1}L_{x}^{\infty}}\leq{\big\|\omega_{0}\big\|}_{X_{0}}.$
This completes the proof. ∎